# Learning to Play Pursuit-Evasion with Dynamic and Sensor Constraints

### v1

@@ IV-C Reward Structure
where $\lambda_{capture}$, $\lambda_{t}$, $\lambda_{d}$ represent the coefficients for capture reward, per-timestep reward and distance reward, respectively. Each episode has two independent reset conditions, the capture condition (pursuer wins) or the timeout (evader wins). $d$ represents the Euclidean distance between agents as described in Section III.
@@ V Simulation Experiments
Figure 3: Learning regime for both agents: Pursuer initially starts out by losing almost every episode, after a few thousand episodes the performance catches up and pursuer starts to dominate. In response, the evader develops its own counter strategies; resulting in an oscillatory reward profile with evader being slightly better on average.
In this section, we analyze our method by evaluating it against pursuer and evader baselines in a series of experiments. We have designed our experiments to answer the following questions: Question 1: How well do trained pursuer and evader agents perform against baseline methods? Question 2: Qualitatively, are there any interesting emerging behaviors? Question 3: How do the model parameters affect agents’ performances?
Figure 4: 250 initial distributions sampled uniformly at random: Initial positions are transformed with respect to the pursuer frame of reference. Colors indicate the normalized time-to-capture. Evader starting points that are farther away from pursuer have higher Time-to-Capture.
@@ V-A Multi-Agent Training
During the training, both agents start from scratch and improve their performance by playing against each other over 15000 episodes. We set $\Delta t=0.1\ s$ and maximum number of timesteps as $\mathcal{T}=500$ with 2-frame stacking and skipping. We spawn 8 parallel vectorized environments to increase training speed with concurrency using the AgileRL Deep Reinforcement Learning Python library MADDPG and environment parallelization [39]. Actor and critic networks are both multilayer perceptrons and comprise of 2 hidden layers of size 128 each and ReLU activations. The workspace $\mathcal{E}$ is of size $(20,20)$ in meters along the x and y-axes and is centered at $(0,0)$. Starting positions of the agents are sampled uniformly at random within $\mathcal{E}$. Reward coefficients are set as $\lambda_{capture}=1000$, $\lambda_{d}=1$, $\lambda_{t}=1$. We use a simple curriculum learning method which starts training with stronger sensor coverage for the pursuer with $V^{p}_{\alpha}=2\pi\ rad$ and $V^{p}_{r}=10\ m$ and linearly decays to the training model parameter over $\frac{1}{4}$ of the number of total training episodes.
@@
For training, pursuer dynamic model parameters are set as follows: rear and front axle distances from the center-of-mass $l_{f},l_{r}=0.15\ m$, steering angle $\theta$ joint limits are ±$\pm 0.34\ rad$, steering angle velocity $\dot{\theta}$ limits are ±$\pm 3.2\ \frac{rad}{s}$. Minimum and maximum longitudinal velocities $v$ are $-1\ \frac{m}{s}$ and $2.5\ \frac{m}{s}$, respectively with ±$\pm 2\ \frac{m}{s^{2}}$ maximum acceleration $\dot{v}$ value in both directions. Sensor parameters are set as $V^{p}_{\alpha}=\frac{\pi}{2}\ rad$ and $V^{p}_{r}=7.5\ m$. Evader dynamic model parameters are: minimum and maximum velocities along the axes are ±$\pm 2\ \frac{m}{s}$ with ±$\pm 9.81\ \frac{m}{s^{2}}$ maximum acceleration $\dot{v}$ value in both directions.
@@
Fig. 3 shows a sample plot of averaged rewards of the agents during training. We observe a learning regime where the pursuer initially starts out losing almost every episode. Eventually the pursuer performance catches up. At that time, the evader develops its own counter strategies; resulting in an oscillatory reward profile. Maximum achievable reward by the pursuer in an individual environment before reset is $\lambda_{capture}=1000$ (instant capture), yet the rewards are accumulated every 500 timesteps for plotting from 8 parallel environments, therefore each parallel environment can reset multiple times (indicating pursuer domination) before the evaluation timestep, resulting a higher cumulative reward. Curriculum learning method described earlier is of higher importance during the early low reward regime for the pursuer; without the curriculum, the pursuer may never be able to encounter the evader and get stuck with a very sparse reward policy, unable to learn.
@@ V-B Performance evaluation against baselines
In our first set of experiments we investigate the performance of our learned pursuer and evader models against baseline algorithms. For multi-agent training scenario, it is not straightforward to find the best performing algorithm by evaluating algorithms based on their reward or similar metrics at the end of an episode due to the oscillatory nature of the training regime. As a more concrete example, a mediocre pursuer may achieve a deceivingly high score after training against a relatively bad evader or vice versa. To mitigate this problem, we create model checkpoints every 250 episodes and evaluate these checkpoints by setting them up against each other (one vs. 64 random samples) after training. The best performing models are selected by evaluating the average success rates and breaking ties with average capture times.
@@
Evader Type Pure Pursuit Learned Pure Pursuit Learned
Random Walk 0.53 ($\sigma$ = 0.35) 0.35 ($\sigma$ = 0.28) 0.77 0.96
Greedy 0.84 ($\sigma$ = 0.28) 0.54 ($\sigma$ = 0.37) 0.30 0.63
Rash 0.81 ($\sigma$ = 0.33) 0.55 ($\sigma$ = 0.39) 0.33 0.62
Learned 0.99 ($\sigma$ = 0.09) 0.69 ($\sigma$ = 0.31) 0.01 0.58
TABLE I: PURSUER CAPTURE TIMES AND RATES
Figure 5: Normalized Time-to-Capture values with respect to the pursuer and evader parameters: Pursuer policies are evaluated against greedy evader policy whereas evader policies are evaluated against learned pursuer policy. Pursuer is able to make good use of its improved sensing capabilities, evader makes rapid gains with velocity until it surpasses the default pursuer velocity.
For learned pursuer and evader strategies, we tested the MADDPG algorithm against baseline pursuer and evader algorithms as well as against each other over 250 episodes. We report the results of the best performing strategies in Table I. The results are normalized between 0 and 1. For the learned pursuer strategy, a lower Time-to-Capture value and a higher capture rate are better. For the learned evader strategy, these values are the opposite. From Table I we see that the learned strategies outperform the baselines in each case and the learned evader strategy performs significantly better against the learned pursuer strategy as compared to the baseline evader strategy. Fig. 4 visualizes these results over 250 initial distributions, uniformly sampled at random from $\mathcal{E}$. The top row visualizes the performance of the baseline pure pursuit where the bottom row uses the learned pursuer strategy. Initial positions are transformed with respect to the pursuer frame of reference (therefore the axes ranges are doubled). Colors indicate the normalized time-to-capture: blue indicates a quicker capture and a stronger pursuer performance; red indicates a slower capture and a stronger evader performance.
V-C Effect of model parameters
Next, we evaluate our approach in the following scenario: pursuer policies are evaluated against greedy evader policy, whereas evader policies are evaluated against learned pursuer policy. We only evaluate by the generalized inference performance, i.e. we do not retrain the agents for the specified parameters and keep rest of the parameters in default settings as described in Section V except for the independent model parameters. Normalized Time-to-Capture values are shown in Fig. 5 with respect to the pursuer and evader parameters. In Figures 5a and 5b it is demonstrated that the learned pursuer policy is able to make good use of its improved sensing capabilities while its ability to handle better forward velocity ends up being in limited Fig. 5c. We attribute this to the relatively limited workspace area, i.e. beyond $4\frac{m}{s}$ it takes less than 5 seconds to go through an axis; higher speeds reducing agility due to braking limits and finite sensor range. The learned evader policy on the other hand does not make good use of its improved sensor range in Fig. 5b but makes rapid gains with velocity in Fig. 5c until it surpasses the default pursuer velocity. Qualitative analysis indicates that the learned evader strategy prefers hiding (close to corners) in favor of outranging pursuer sensor, only moving away if the pursuer decides to check the corner the evader is currently in. This, however comes with a significant drawback, i.e. the evader corners itself even with clear dynamic superiority and gets captured. This indicates that the qualitative nature of learned strategies are strongly correlated with model parameters.
VI Real-World Results
For the real world experiments, a learned pursuer policy was deployed against a human controlled evader. The reinforcement learning model for the pursuer was deployed on an F1TENTH autonomous vehicle [6] with the identified system parameters from our previous work [40]. The human-controlled evader was a smaller and more nimble car-like autonomous vehicle Nvidia JetRacer [7]. Due to real world physical and system constraints, the reinforcement learning model state-space representation was slightly modified to use linear velocity and steering angle commands (as opposed to their first order derivatives), full observability was assumed (due to area being too small to properly make use of sensor footprint) and steering angles are removed from the observation state (since they are actions). A Gaussian noise with 0 mean and 5% standard deviation was added to the normalized (between -1 and 1) state and action vectors for robustness. For 6-DoF real-world state estimation, Phasespace X2E LED motion capture markers were attached to both vehicles and finite differencing was used to calculate velocities at 10 Hz frequency. ROS2 was used as the middleware for interfacing between sensors, agents and the reinforcement learning models [41]. We demonstrate trained reinforcement learning models are directly transferable to the real system and present two sample pursuit evasion sample trajectories in Fig. 6, for further details on our physical experimental results, please refer to the submitted video attachment. These experiments successfully demonstrate that the learned strategy can be executed on real systems.
Figure 6: Real system trajectories: We demonstrate trained reinforcement learning models are directly transferable. Lighter colors have earlier timestamps, darker colors are later. For further details on our physical experimental results, please refer to the submitted video attachment.
VII Conclusion
@@
In this paper, we explored a pursuit-evasion game where the pursuer, a car with five state variables, engages with either a similar car or a point mass evader with constraints on the sensing capabilities for both agents. We formulated the game as a zero-sum Partially Observable Stochastic Game (POSG) and developed a curriculum for the MADDPG algorithm to simultaneously obtain the pursuer and evader policies. By training the players against each other, our approach sidesteps the necessity for pretrained or expert-designed evader policies. We also demonstrated that simulation-trained policies are directly transferable to real autonomous agents. Through this exploration, we show the potential of multi-agent reinforcement learning in navigating complex pursuit-evasion scenarios within dynamic environments.
@@
There are multiple avenues for future research: in our current formulation, we did not model uncertainty in the players own states. Preliminary experiments indicate that the learned strategies are robust to small errors in state. As the arena gets bigger and if the players do not have perfect global sensors, the uncertainty in state estimation may need to be addressed explicitly. Furthermore, if the game takes place in complex environments, the search component of the game becomes prominent. It is unclear if current RL methods will be able to handle such scenarios in which there will be long sequences with no reward. We will explore these exciting, yet challenging, issues in our future work.

### v2

@@ IV-D Reward Structure
where $\lambda_{capture}$, $\lambda_{t}$, $\lambda_{d}$ represent the coefficients for capture reward, per-timestep reward and distance reward, respectively. Each episode has two independent reset conditions, the capture condition (pursuer wins) or the timeout (evader wins). $d$ represents the Euclidean distance between agents as described in Section III. Per-timestep rewards encourage the pursuer to capture as quickly as possible while distance rewards provide a denser reward structure to help the training procedure.
Figure 4: BiMDN architecture: BiMDN uses 1 layer of bidirectional LSTM, with 128 hidden units followed by a fully connected layer of 32 output features. The network thereafter splits into 3 heads with a fully connected layer in each head before the output layers. Outputs’ activation functions are softmax, Exponential Linear Unit (offset with +1 to enforce nonnegativity) for cluster weights $\Pi^{i}_{t}$ and standard deviations $\Sigma^{i}_{t}$, respectively. Gaussian cluster centers $M^{i}_{t}$ are directly predicted by fully connected layers. Fully connected layers are represented in red (not the actual number of parameters) and number of output clusters is empirically set at 3.
@@ V Simulation Experiments
In this section, we analyze our method by evaluating it against pursuer and evader baselines in a series of experiments. We have designed our experiments to answer the following questions: Question 1: How well do trained pursuer and evader agents perform against baseline methods? Question 2: Does BiMDN improve model performance over existing MARL methods?
@@ V-A Multi-Agent Training
During the training of our method, both agents start from scratch and improve their performance by playing against each other over 20000 episodes. We set $\Delta t=0.1\ s$ and maximum number of timesteps as $\mathcal{T}=400$ with 2-frame stacking and skipping. We spawn 8 parallel vectorized environments using the AgileRL Deep Reinforcement Learning Library [60]. Actor and critic networks are both multilayer perceptrons (MLP) and comprise of 2 hidden layers of size 128 each and ReLU activations. We set replay buffer size at 100000 with 512 batch size, learning rate at 0.0005 for both actor and critic networks with Adam [61] optimizer, exploration noise as Gaussian noise with 0.1 standard deviation and 0 mean. Discount factor ($\gamma$) and soft target update parameter ($\tau$) are set at 0.99 and 0.005, respectively. In our simulation experiments BiMDNs use 1 layer of bidirectional LSTMs, which contain 2 seperate LSTMs for forward and backward passes with 128 hidden units each followed by a fully connected layer of 32 output features. The network thereafter splits into 3 heads with a fully connected layer in each head before the output layers. For cluster weights $\Pi^{i}_{t}$, the number of output features is equal to the number of cluster centers in the corresponding layer, other heads have twice the number of output features, representing 2D coordinates. Outputs’ activation functions are softmax, Exponential Linear Unit (offset with +1 to enforce nonnegativity) for cluster weights $\Pi^{i}_{t}$ and standard deviations $\Sigma^{i}_{t}$, respectively. Cluster centers $M^{i}_{t}$ are directly predicted by the fully connected layers and number of clusters is empirically set at 3. BiMDN architecture is presented in Fig. 4. 200 previous observations are used for generating the belief states, downsampled by the factor of 10, corresponding to $1\ Hz$. BiMDNs have their own buffers of size 25000 and have a warm-up period of 10000 samples. After that, belief networks are updated every 10 timesteps with samples collected from the corresponding buffers; in batches of 256 samples with Adam optimizer and learning rate of $0.002$. The workspace $\mathcal{E}$ is of size $(16,16)$ in meters along the x and y-axes and is centered at $(0,0)$. Starting positions of the agents are sampled uniformly at random within $\mathcal{E}$. Reward coefficients are set as $\lambda_{capture}=1000$, $\lambda_{d}=1$, $\lambda_{t}=1$. We use a simple curriculum learning method which starts training with stronger sensor coverage for the pursuer with $V^{p}_{\alpha}=2\pi\ rad$ and linearly decays to the training model parameter over $\frac{3}{10}$ of the number of total training episodes. We also start evader velocity at $0\ \frac{m}{s}$ and increase linearly over the same curriculum period until it reaches its defined value. We empirically find that having a curriculum strategy greatly helps the pursuer at the initial training episodes as exploration is guided poorly by random noise, resulting in a very sparse reward scheme.
@@
For training, pursuer dynamic model parameters are set as follows: rear and front axle distances from the center-of-mass $l_{f},l_{r}=0.15\ m$, steering angle $\theta$ joint limits are ±$\pm 0.34\ rad$, steering angle velocity $\dot{\theta}$ limits are ±$\pm 3.2\ \frac{rad}{s}$. Minimum and maximum longitudinal velocities $v$ are $-1\ \frac{m}{s}$ and $2.5\ \frac{m}{s}$, respectively with ±$\pm 2\ \frac{m}{s^{2}}$ maximum acceleration $\dot{v}$ value in both directions. Sensor parameters are set as $V^{p}_{\alpha}=\frac{2\pi}{3}\ rad$ and $V^{p}_{r}=6\ m$. Evader dynamic model parameters are: minimum and maximum velocities along the axes are ±$\pm 1.5\ \frac{m}{s}$ with ±$\pm 9.81\ \frac{m}{s^{2}}$ maximum acceleration $\dot{v}$ value in both directions.
@@
In Fig. 5 we observe a learning regime where the pursuer initially starts out losing almost every episode. Eventually the pursuer performance catches up with the help of the curriculum strategy. Maximum achievable reward by the pursuer in an individual environment before reset is $\lambda_{capture}=1000$ (instant capture), yet the rewards are accumulated every 500 timesteps for plotting from 8 parallel environments, therefore each parallel environment can reset multiple times (indicating pursuer domination) before the evaluation timestep, resulting a higher cumulative reward. We propose two variants derived from our BiMDN method, the first one uses weighted mean of Gaussians as the belief state vector during inference time, which represents a pure strategy (Ours), whereas the second variant samples 16 points from the mixture, representing a mixed strategy (Ours-Mixed).
Figure 5: Learning regime for both agents: Pursuer initially starts out by losing almost every episode, after a few hundred episodes the performance catches up and pursuer starts to dominate with the help of curriculum strategy. In response, evader develops its own counter strategies.
@@ V-B Performance evaluation against baselines
In our first set of experiments we investigate the performance of our learned pursuer and evader models against baseline algorithms. For multi-agent training scenario, it is not straightforward to find the best performing algorithm by evaluating algorithms based on total reward or similar metrics at the end of an episode due to the oscillatory nature of the training regime. As a more concrete example, a mediocre pursuer may achieve a deceivingly high score after training against a relatively bad evader or vice versa. To mitigate this problem, we create model checkpoints every 250 episodes and evaluate these checkpoints by setting them up against each other (one vs. 32 random samples) after training. The best performing models are selected by evaluating on average capture rates.
Evader Type Pure Pursuit MADDPG UKF[58] Image[31] Ours Ours-Mixed
Random Walk 0.59 ($\sigma$=0.36) / 0.67 0.23 ($\sigma$=0.29) / 0.92 0.27 ($\sigma$=0.22) / 0.98 0.32 ($\sigma$=0.27) / 0.94 0.27 ($\sigma$=0.20) / 0.98 0.29 ($\sigma$=0.22) / 0.98
Greedy 0.79 ($\sigma$=0.33) / 0.34 0.44 ($\sigma$=0.30) / 0.83 0.31 ($\sigma$=0.21) / 0.94 0.36 ($\sigma$=0.25) / 0.93 0.30 ($\sigma$=0.16) / 0.98 0.31 ($\sigma$=0.15) / 0.99
Rash [57] 0.74 ($\sigma$=0.37) / 0.39 0.39 ($\sigma$=0.32) / 0.82 0.37 ($\sigma$=0.32) / 0.83 0.40 ($\sigma$=0.34) / 0.82 0.29 ($\sigma$=0.23) / 0.94 0.32 ($\sigma$=0.20) / 0.99
MADDPG 0.86 ($\sigma$=0.28) / 0.27 0.52 ($\sigma$=0.36) / 0.69 0.42 ($\sigma$=0.32) / 0.79 0.45 ($\sigma$=0.31) / 0.84 0.34 ($\sigma$=0.22) / 0.96 0.34 ($\sigma$=0.21) / 0.98
UKF[58] 0.95 ($\sigma$=0.19) / 0.09 0.44 ($\sigma$=0.30) / 0.84 0.49 ($\sigma$=0.36) / 0.70 0.51 ($\sigma$=0.34) / 0.75 0.36 ($\sigma$=0.27) / 0.89 0.36 ($\sigma$=0.20) / 0.97
Image[31] 0.95 ($\sigma$=0.19) / 0.09 0.54 ($\sigma$=0.36) / 0.68 0.46 ($\sigma$=0.32) / 0.82 0.54 ($\sigma$=0.34) / 0.75 0.34 ($\sigma$=0.22) / 0.95 0.36 ($\sigma$=0.22) / 0.97
Ours 0.92 ($\sigma$=0.22) / 0.15 0.43 ($\sigma$=0.30) / 0.83 0.52 ($\sigma$=0.35) / 0.70 0.54 ($\sigma$=0.33) / 0.73 0.36 ($\sigma$=0.26) / 0.89 0.35 ($\sigma$=0.18) / 0.99
Ours-Mixed 0.91 ($\sigma$=0.22) / 0.17 0.51 ($\sigma$=0.35) / 0.69 0.47 ($\sigma$=0.31) / 0.86 0.50 ($\sigma$=0.34) / 0.74 0.30 ($\sigma$=0.17) / 0.99 0.38 ($\sigma$=0.19) / 0.99
TABLE I: PURSUER CAPTURE TIMES, STANDARD DEVIATIONS / RATES
For learned pursuer and evader strategies, we tested our method against baseline pursuer and evader algorithms as well as against each other over 1500 episodes, using 3 different trained models with 3 random seeds, 500 episodes each. We report the results of the best performing strategies in Table I. The results are normalized between 0 and 1 to represent rates. For any pursuer strategy, a lower capture time and a higher capture rate are better. For any evader strategy, these values are the opposite. From Table I we see that our proposed methods both surpass existing baselines, which indicates that BiMDN improves capture performance in partially observable games. Kalman Filter and image representations also improve over vanilla MADDPG against heuristic evader baselines; however their advantages diminish once they play against learning-based policies, as their pursuer strategies end up being exploitable. The advantage is less pronounced for our trained evader strategies as they struggle to outperform the baseline strategies; showing that belief states are more useful for improving pursuer strategies, although MARL methods are still able to train relatively strong evader policies. To test our method for statistical significance, we use Two Proportion Z-Test over the capture rates’ distributions and represent the strongest policies using bold font. Two methods having bold font in the same row means that the p-value is over 0.05 between two distributions.
V-C Performance evaluation against Hamilton-Jacobi solvers
In this section we used OptimizedDP [19] to solve a differential game. Such solvers have a large memory footprint, therefore we reformulate the problem to be a differential game between two homogeneous point-mass agents, reduce the environment size and allowable game duration to achieve computational feasibility. The workspace $\mathcal{E}$ is of size $(10,10)$ in meters; maximum number of timesteps is $\mathcal{T}=100$ without frame stacking or skipping. Sensor ranges are $3.75\ m$ and maximum velocities along both axes are increased to $2.5\ \frac{m}{s}$. Our method is trained for 20000 episodes. We follow the same experimental procedure as Section V-B and report the results in Table II. Since optimal control policies are generated for fully observable cases we use BiMDN weighted mean from the Gaussian mixture when the opponent is not observable and compare against the pure strategy accordingly. From Table II we see that our pursuer strategies are significantly better than the optimal control policies. Our pursuer method improves over existing baselines by a statistically significant margin. The comparative advantage provided by our method to the evader policies is less pronounced with the optimal control pursuer policy being slightly better $(p=0.03)$ against our evader method in terms of capture rate, with the help of the state estimation provided by BiMDN. Results indicate that our proposed method is also competitive with baselines from the reachability literature for simpler problems that are computationally feasible.
VI Real-World Experiments
For the real world experiments, a learned pursuer policy was deployed against a human controlled evader. Pursuer policy was directly implemented on an F1TENTH autonomous vehicle [9] with the identified system parameters from our previous work [62]. The human-controlled evader was a more nimble autonomous vehicle Nvidia JetRacer [10]. Sensor ranges were set at $5\ m$ and pursuer sensor field of view angle was set at $\frac{3}{2}\pi\ rad$. The real-world sensor coverage provided by the F1TENTH LiDAR goes up to $20\ m$ at $\frac{3}{2}\pi\ rad$, however we artificially restrict the simulated sensor range at $5\ m$ as our testing environment is significantly smaller at $(10,5.5)$ in meters along the x and y-axes. For 6-DoF real-world state estimation, Phasespace X2E LED motion capture markers were attached to both vehicles and finite differencing was used to calculate velocities at $10\ Hz$. ROS2 was used as the middleware [63]. For further details on our physical experimental results, please refer to the submitted video attachment. These experiments successfully demonstrate that the learned strategy can be executed on real systems.
VII Conclusion
@@
Evader Type Optimal Control Ours Optimal Control Ours
Optimal Control [19] 0.62 ($\sigma$ = 0.40) 0.46 ($\sigma$ = 0.32) 0.47 0.83
Ours 0.62 ($\sigma$ = 0.40) 0.47 ($\sigma$ = 0.32) 0.50 0.80
TABLE II: POINT-MASS CAPTURE TIMES AND RATES
@@
In this paper, we explored a pursuit-evasion game where the pursuer, a car with five state variables, engages with either a similar car or a point mass evader with constraints on the sensing capabilities for both agents. We formulated the game as a zero-sum Partially Observable Stochastic Game (POSG) and developed a bidirectional recurrent mixture model named BiMDN for complementing the MADDPG algorithm to simultaneously obtain the pursuer and evader policies. By training the players against each other, our approach sidesteps the necessity for pretrained or expert-designed evader policies. We also demonstrated that policies obtained in simulation are directly transferable to real autonomous agents and our method is able to handle homogeneous as well as heterogeneous cases in terms of dynamics governing the agents, where existing level-set methods are infeasible. Through this exploration, we show the potential of multi-agent reinforcement learning in navigating complex pursuit-evasion scenarios within dynamic environments.
@@
There are multiple avenues for future research: in our current formulation, we did not model uncertainty in the players own states. Preliminary experiments indicate that the learned strategies are robust to small errors in state. As the arena gets bigger and if the players do not have perfect global sensors, the uncertainty in state estimation may need to be addressed explicitly. Furthermore, if the game takes place in complex environments, the search component of the game becomes prominent. It is unclear if current methods will be able to handle such scenarios in which there will be long sequences with no reward. We will explore these exciting, yet challenging, issues in our future work.